# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 2. Background
Given $\epsilon>0$, and $\delta\in(0,1)$. A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all pairs of neighboring databases $D,D^{\prime}$, and any possible output $o\subset Range(\mathcal{M})$, the following holds:
$\textup{Pr}\left[\mathcal{M}(D)\in o\right]\leq e^{\epsilon}\textup{Pr}\left[\mathcal{M}(D^{\prime})\in o\right]+\delta$
3. Privacy Model
In this section, we formally define the privacy definitions of SPECIAL and explain the corresponding semantics. Similar to previous work on DPSCA (Wang et al., 2022; Zhang et al., 2023), we follow computational DP (SIM-CDP) (Mironov et al., 2009) framework for privacy model formulation.
Definition 3.1 (Secure protocol with DP leakage).
Given a party (owner) $P$ with private data $D$ and a randomized mechanism $\mathsf{Lkg}(D)$, referred to as the leakage profile of $P$’s data. Consider $\Pi$ a secure query processing protocol applied to $D$. The protocol $\Pi$ is said to be secure with DP leakage if, for any honest party $P$ with data $D$ and any probabilistic polynomial time (p.p.t.) adversary $\mathcal{A}$:
  • $\mathsf{Lkg}(D)$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
@@ 3. Privacy Model
  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and the output of $\mathsf{Lkg}(D)$ that satisfies:
@@
    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(F(\mathcal{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
@@
where $\mathsf{VIEW}^{\Pi}$, and $\mathsf{VIEW}^{\mathcal{S}}$ denotes the adversary’s view against the protocol execution and the simulator’s outputs, respectively. And $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
Simply put, Definition 3.1 specifies that as long as there is at least one honest owner, the knowledge any p.p.t. adversary can obtain about this owner’s data from observing $\Pi$ execution transcripts is bounded to what can be inferred from the outputs of an $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$. By default, we define $\mathsf{Lkg}$ within $(\epsilon,\delta)$-tuple level DP (Dwork et al., 2010), that’s said primary privacy objective is to protect individual tuples. However, due to group-privacy properties (Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015) of DP, privacy can extend across multiple tuples and to user-level, as long as a user owns a limited number of tuples. Throughout this paper, our focus remains on developing algorithms that guarantee tuple-level privacy, while upholding all stated privacy guarantees, albeit with potentially different privacy parameters.
Theorem 3.2.
The privacy of SPECIAL adheres to Definition 3.1
Proof.
Due to space concerns, we defer the complete privacy proof of SPECIAL to Appendix A. ∎
@@ 4. SPECIAL Synopses
SPECIAL maintains independent DP synopses for all outsourced data. Each synopsis holds noisy statistics that reflect the corresponding data’s distributional traits. As per our core design idea, these statistics will be used to dynamically optimize query processing—for example, by estimating tight compaction bounds or enabling fast access to eliminate unnecessary tuples, which significantly compacts intermediate query sizes, all without any extra privacy cost.
Challenges. A major design challenge is selecting proper synopses to achieve our design goals. This involves decisions such as attribute selection, combinations thereof, and the types of statistics used (e.g., histograms, min/max/avg, frequency counts, etc.). Using too many attributes or leveraging complex (high-dimensional) combinations can introduce significant noise (Zhang et al., 2021) and increase storage costs. Another challenge involves privately releasing synopses to aid in lossless processing, such as result compaction and private indexing. Ensuring lossless compaction requires consistently overestimating the output sizes of subplans. Supporting lossless indexes is even more complex: for a given indexed range $D[\mathsf{lo},\mathsf{hi}]$, we need to overestimate the upper index $\mathsf{hi}$ and underestimate the lower index $\mathsf{lo}$. Unfortunately, traditional DP methods, which typically use symmetric additive noises, struggle to ensure these conditions.
Key ideas. We observe that joins are more resource-intensive than any other operations. For instance a $k$-way oblivious join can incur a $O(n^{k})$ complexity if without optimisations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017). Hence, joins require prioritized acceleration. Moreover, synopses available for frequently queried filter attributes on base relations are vital (Wang et al., 2022) as well. These statistics not only help to compact input data at an early stage before heavy joins, but also enable fast indexing (§ 5.2) for quick data access. Hence, to address the first challenge, our key idea is to focus on commonly queried join and base relation filter attributes, as well as their low-dimensional combinations.
To tackle the second challenge, our key idea is twofold. First, to ensure lossless result compaction of filters and indexing, we employ one-sided DP noise to generate private histograms that either overestimate or underestimate attribute distributions consistently. Second, to achieve lossless compaction of join outputs, we integrate noisy max frequency moments (MF) into the synopses. This integration enables us to extend upon advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes without losing data (§ 5.2).
@@ 4.1. Synopses generation
Steered by our key ideas, we now elaborate on our synopsis generation details. The general process is outlined as follows: Initially, the SPECIAL servers selects appropriate attributes to generate private synopses, which are then distributed to owners. The owners create the synopses using a DP mechanism and upload them to SPECIAL servers. Finally, the server applies post-processing to these synopses to aid in query processing.
Attributes selection (servers). The first step is to identify a smaller set of attributes for deriving synopses. In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run of SPECIAL or annotated by the administrator. Note that this process do not incur any private data thus is leakage-free. The servers first identify representative attribute pairs, $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$, for each private relation $D\in\mathcal{D}$ via $Q_{\mathsf{R}}$. The designated pairs include: (i) 2-way attribute pairs, which correspond to frequently queried filter-join key combinations; (ii) frequently queried individual attributes not covered by these pairs. By default, each $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})$ contains two valid attributes (case i), but either $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ may be empty (case ii).
Synopses release (owners). Next, servers pushes the identified pairs to each owner, and subsequently, the owner independently dispatches a private synopsis and returns it to servers. In what follows, we focus on the DP synopses generation mechanism evaluated by each owner, Algorithm 1 illustrates the overall flow.
Algorithm 1 DP synopsis generation $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
@@
8$\triangleright$ adding independently sampled noise to every bin of $\mathbf{h}^{+}$, $\mathbf{h}^{-}$
@@
12$\triangleright$ assuming $\mathbf{h}$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$
@@
Generally, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$) before initiating the synopsis generation. For each $\mathsf{pair}_{k}$, the owner first creates a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$. By default, we assume there exist global parameters, i.e. bin sizes of each attribute, which guide all owners to partition attributes consistently. Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$, and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$ by adding independently sampled one-sided Laplace noises to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$.
@@
$\mathsf{Lap}^{+}(\epsilon,\delta)=\max(0,z)$ (resp. $\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $[0,\infty)$ (resp. $(-\infty,0]$) if $z$ is drawn from a distribution with the following density function
@@
The one-sided Laplace noise can be strictly positive ($\mathsf{Lap}^{+}$) or negative ($\mathsf{Lap}^{-}$). This ensures that $\mathbf{h}^{+}$ always overestimates the actual histogram, while $\mathbf{h}^{-}$ consistently underestimates it. For the ease of description, we refer $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ as the bounding histogram of attribute $\mathsf{pair}_{k}$, with $\mathbf{h}^{+}$, and $\mathbf{h}^{-}$ to be the upper and lower histograms, respectively.
We continue with the details of the (noisy) MFs generation. Assume both $A_{\mathsf{ft}},A_{\mathsf{j}}\neq\emptyset$, and the $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$. The mechanism $\mathcal{M}_{\mathtt{synop}}$ generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{\ell\geq 1}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ is an independent MF statistic for $A_{\mathsf{j}}$ over specific filtered data on $A_{\mathsf{ft}}$, such that
@@
here $\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}$ is a group-by-count operation over $A_{\mathsf{j}}$, and $\widehat{\max}_{\epsilon}$ is a report noisy max mechanism (Dwork et al., 2014). It first adds i.i.d. noise from the exponential distribution $\mathsf{Exp}(\frac{2}{\epsilon})$ to each grouped count, then outputs the largest noisy count. We stress that $\mathcal{M}_{\mathtt{synop}}$ will not generate noisy MFs for non-join key attributes, and when $A_{\mathsf{ft}}$ is empty, a global MF will be generated instead of MF tables (Alg 1:10). Moreover, since SPECIAL enables owners to label attributes as unique-valued, if $A_{\mathsf{j}}$ is known to be unique-valued, then $\smash{\widehat{\mathtt{mf}}(A_{\mathsf{j}},\cdot)}$ is always 1. Nevertheless, as exponential noises are non-negative, thus $\widehat{\mathtt{mf}}\geq\mathtt{mf}$ holds for all cases. Given the discussion so far, we summarize SPECIAL synopses as:
Definition 4.2 (SPECIAL synopses).
Given representative workload $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ to be the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
  • $\mathsf{pair}_{k}$ is a frequently queried attribute (pair) over $D$ found in $Q_{\mathsf{R}}$.
  • $\mathbf{H}$ is the private bounding histogram for $\mathsf{pair}_{k}$ over $D$.
  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs that corresponds to a group of tuples in $\mathbf{h}$, categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
@@
Given $|\mathsf{pair}|=c$, $\epsilon,\delta>0$, the synopsis generation (Algorithm 1) is $(\hat{\epsilon},\hat{\delta})$-DP where $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$, and $\hat{\delta}=(c+1)\delta$
For space concern, we move the complete proof to Appendix A. In a sketch, adding $\mathsf{Lap}^{+}$ (or $\mathsf{Lap}^{-}$) to a single bin is $(\epsilon,\delta)$-DP. By parallel and sequential composition, generating $\mathbf{{H}}$ is $(2\epsilon,2\delta)$-DP. Moreover, each report noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP. In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,\delta)$-DP. Given there are in total $c$ such pairs, and thus the total privacy loss is subject to $c$-fold advanced composition (Dwork et al., 2014).
Synopsis transformations. We say that one can perform transformations on released synopses without incurring extra privacy loss, per the post-processing theorem of DP (Dwork et al., 2014). Now, we briefly outline key synopsis transformations relevant to the SPECIAL design. First, given any (2d) bounding histogram $\mathbf{H}(\mathsf{pair},D)$ with both $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}$ are non-empty, one can derive the (1d) bounding histograms, i.e. $\mathbf{H}(A_{\mathsf{j}},D)$ and $\mathbf{H}(A_{\mathsf{ft}},D)$, for any single attribute $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ by marginal sums $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over $A_{\mathsf{j}}$ or $A_{\mathsf{ft}}$, respectively. This enables the creation of statistics on individual attributes, even when $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are not included as a standalone synopsis attribute. Moreover, it’s possible to derive relevant join key statistics even following a selection on the base relation. For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with $vals$. Beyond bounding histograms, join key MFs over pre-filtered data can also be computed by

### v2

@@ 2. Background
Given $\epsilon>0$, and $\delta\in(0,1)$. A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all ${D}\sim{D^{\prime}}$ pairs, and any possible output $o\subset Range(\mathcal{M})$, the following holds: Pr[M(D) o] e^ϵ Pr[M(D’) o] + δ
[…]
Oblivious (relational) operators. Oblivious operators are data-independent MPC protocols that implement the same functionalities as their plaintext database counterparts (e.g., filter and join). Data-independent execution requires that the control flow and memory access patterns of a function are indistinguishable given different inputs of the same size, and typically requires costly computation. For example, a linear scan is required to fulfill oblivious filtering (Zheng et al., 2017), and join requires nested-loop over the two inputs (Eskandarian and Zaharia, 2017). The output sizes of such operators are usually padded with dummy tuples to the worst case: $N$ rows for filters and $N^{2}$ for joins, given size $N$ inputs. The dummy tuples will not affect the query result but can significantly impact the performance (Zheng et al., 2017; Eskandarian and Zaharia, 2017). To enhance efficiency while maintaining strong privacy, DPSCAs introduce a new type of oblivious operators (Bater et al., 2018; Qin et al., 2022; Qiu et al., 2023; Wang et al., 2022). These operators typically involve two steps: Compute and Compact. The Compute step is fully oblivious, while the Compact resizes the output, often by obliviously sorting valid tuples to the front and trimming the output to a noisy DP size (the true size plus DP noise). While this approach can significantly reduce the computation cost and query sizes, it may lead to lossy query processing if the DP size is smaller than the true size (e.g., negative DP noise), as valid tuples could be excluded during the compaction (Wang et al., 2021, 2022; Zhang et al., 2023). We emphasize that when DP sizes exceed true sizes, there is no accuracy loss as it only includes extra dummy data that do not impact accuracy (Zheng et al., 2017; Eskandarian and Zaharia, 2017).
Private indexes. In conventional databases, indexes are powerful data structures that map attribute values to positions in a sorted array, allowing a predicate selection to quickly access the desired data via index lookup without the need for full table scan. However, traditional indexes are unsuitable in SCAs due to their data-dependent nature, which can easily lead to privacy breaches. To address this, recent research has proposed DP indexes (Roy Chowdhury et al., 2020), where the mapping of attribute values to their positions is intentionally distorted with DP noise. To process queries, the system first pre-fetch a small set of data using DP indexes, followed by oblivious selection. This effectively avoids full table oblivious scan and sorting-based result compaction. However, the uncertainty inherent in DP indexes can lead to the loss of valid tuples. For example, in the DP index of (Roy Chowdhury et al., 2020), a true index range of positions $[10,20]$ might be distorted into positions $[12,17]$, causing data at position $10,11,18,19$ and $20$ to be missed. Nevertheless, if DP indexes overestimate the range, subsequent oblivious selection can losslessly identify all valid tuples.
3. System and Privacy Model
In general, SPECIAL follows a standard server-aided MPC (Kamara et al., 2011) model, involving (i) a set of mutually distrustful data owners $P_{1},...,P_{n}$}, (ii) two non-colluding servers $S_{0}$ and $S_{1}$, and (iii) a trusted analyst. We assume an admissible adversary (Mohassel and Zhang, 2017) $\mathcal{A}$, capable of corrupting n1n-1 out of $n$ clients and at most one of the two servers. An instance of such adversary can be a malicious server that creates Sybil owners to form a malicious collation, attempting to steal sensitive information from an honest owner. Additionally, $\mathcal{A}$ is considered honest-but-curious, meaning it follows the protocol without deviation but may try to infer information from observed protocol transcripts, such as randomness, memory access patterns, and communication messages. The combination of these information is referred to as the view of $\mathcal{A}$. We also assume $\mathcal{A}$ is computationally bounded as a probabilistic polynomial time (p.p.t.) adversary, which is a standard requirement in MPC protocols to ensure that adversaries cannot break cryptographic primitives. This threat model is consistent with prior SCA designs (Bater et al., 2017, 2018; Wang et al., 2022; Mohassel and Zhang, 2017; Wang et al., 2021). Given this setup, we design SPECIAL to satisfy the following:
Definition 3.1 (MPC protocol with DP leakage).
Given a set of parties (owners) $P_{i}$ with private data $D_{i}$ and a secure query protocol $\Pi$ that applies over $\mathcal{D}=\{D_{1},D_{2},\cdots\}$. We define a randomized mechanism $\mathsf{Lkg}(\mathcal{D})=\{\mathsf{Lkg}(D_{1}),\mathsf{Lkg}(D_{2}),\cdots\}$ as the leakage profile, consisting of the control flow and access patterns of running $\Pi$ over $\mathcal{D}$. The protocol $\Pi$ is said to be secure with DP leakage if, for the subset of uncorrupted parties with data $\mathbf{D}\subseteq\mathcal{D}$, leakage profile $\mathsf{Lkg}(\mathbf{D})\subseteq\mathsf{Lkg}(\mathcal{D})$, and any p.p.t. adversary $\mathcal{A}$:
  • $\mathsf{Lkg}(\mathbf{D})$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
@@ 3. System and Privacy Model
  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and $\mathsf{Lkg}(\mathbf{D})$ that satisfies:
@@
    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(\mathsf{Lkg}(\mathbf{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
@@
where $\mathsf{VIEW}^{\Pi}$ is $\mathcal{A}$’s view in $\Pi$’s execution and $\mathsf{VIEW}^{\mathcal{S}}$ is a simulated view produced by $\mathcal{S}$ using $\mathsf{Lkg}$; $\mathsf{pp}$ denotes all public parameters, and $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
Simply put, Definition 3.1 requires that the knowledge any p.p.t. adversary adversary can gain about each individual tuple of an honest owner, by observing the protocol execution, is bounded to what can be inferred from the outputs of the $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$. We stress that this notion focuses on DP at the event (tuple) level without loss of generality. Due to the group-privacy properties of DP (Dwork et al., 2010; Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015; Vadhan, 2017), event-level DP can be extended to user-level DP. For instance, in a logical database $\mathcal{D}$ where any single user owns at most $l$ tuples, if a protocol satisfies $(\epsilon,\delta)$ event-level DP, it also satisfies $(l\epsilon,le^{(l-1)\epsilon}\delta)$ user-level DP. Moreover, we say that SPECIAL can be relaxed to employ a weaker corruption model, such as requiring a supermajority of owners and servers to remain uncorrupted, to enhance efficiency (Liagouris et al., 2023; Tan et al., 2021). This adjustment does not change the privacy guarantee outlined in Definition 3.1, but it does affect the security assumptions. Under the relaxed corruption model, Definition 3.1 is only satisfied when at least two-thirds of the parties remain uncorrupted. Due to space concerns, we defer the complete privacy proof of SPECIAL to our full version (Chenghong Wang, 2024).
@@ 4. SPECIAL Synopses
We now discuss the details of private synopses used in SPECIAL, while in later sections we will show how they accelerate query processing (§ 5) and aid in query planning (§ 6).
Challenges. We reiterate the main challenges in generating private synopses for a relation include: (C-1) selecting a set of attribute combinations that enable functional and efficient query processing; (C-2) ensuring that the subsequent query processing based on the private synopses is lossless.
Key ideas. Given a SPJA query, join operations typically need prioritized acceleration, as they are more resource-intensive than other operations. A $k$-way join can have $O(n^{k})$ complexity without optimizations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017). Additionally, synopses for frequently queried filter attributes play an important role in efficient query processing, as they enable fast indexing (§ 5.2) and effective filtering of unnecessary data before heavy joins. As such, our first key idea is to focus on histograms-based synopses that cover frequently queried join and filter attributes. To minimize noise, we focus on low-dimensional synopses: only 1D or 2D histograms.
To address the second challenge, our approach combines two strategies. First, to support accurate indexing and filtering, we use one-sided DP noise to generate DP histograms that consistently overestimate or underestimate attribute distributions. We will show later that such special histograms allow lossless filtering and indexing that reliably overestimate true filter sizes and indexing ranges (§ 5.2). Second, for lossless join output compaction, we incorporate noisy max frequency moments (MF) into the synopses. MF allows us to build on advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join sizes without data loss (§ 5.2).
@@ 4.1. Synopses generation
We now elaborate on the details of synopsis generation, which mainly contains two phases: (i) Attributes selection, where the SPECIAL servers select appropriate attributes for the generation, which are then distributed to owners; (ii) Local synopses release, where the owners create corresponding synopses using a DP mechanism and upload them to SPECIAL servers.
Attributes selection (servers). The first step is to identify a set of attributes for deriving synopses. In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run or annotated by the administrator. Note that the representative workload does not involve any private data and thus is leakage-free. The servers first identify representative attribute pairs, $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$, for each private relation $D\in\mathcal{D}$ via $Q_{\mathsf{R}}$. The designated pairs include: (i) 2-way attribute pairs, which correspond to frequently queried filter-join key combinations; (ii) frequently queried individual attributes not covered by these pairs. By default, each $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})$ contains two valid attributes (case i), but either $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ may be empty (case ii).
Synopses release (owners). Next, servers pushes the identified pairs to owners, and subsequently, the owners independently dispatche private synopses and return them to servers. We now focus on the DP synopses generation mechanism run by each owner, Algorithm 1 illustrates the workflow.
Algorithm 1 DP synopsis gen $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
@@
[…]
@@
[…]
@@
In general, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$). Algorithm 1 produces synopses formalized as :
Definition 4.1 (SPECIAL synopses).
Given $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ is the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
  • $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})\in Q_{\mathsf{R}}$ is a frequently queried attribute pair.
  • $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ is the DP bounding histogram for $\mathsf{pair}_{k}$, where $\mathbf{h}^{+}$ (resp. $\mathbf{h}^{-}$) is a DP histogram that overestimate (resp. underestimate) the true histogram of $\mathsf{pair}_{k}$.
  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
We now present a detailed explanation of generating these synopsis structures, starting with the private bounding histograms (Alg 1, lines 2:5). Specifically, for each $\mathsf{pair}_{k}$, each owner first constructs a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$. By default, we assume there exists global parameters (e.g., bin sizes) for each attribute, that ensure consistent data partitioning among all owners. Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$ and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$, by adding independently sampled one-sided Laplace noise (Definition 4.2) to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$. This guarantees that $\mathbf{h}^{+}$ always overestimates the true histogram, while $\mathbf{h}^{-}$ consistently underestimates it.
@@
$\mathsf{Lap}^{+}(\epsilon,\delta)=\max(0,z)$ (resp. $\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $0,\infty)$ (resp. $(-\infty,0]$) if $z$ is drawn from a distribution with the following density function
@@
Next, we detail the generation of (noisy) join key MFs (Alg 1 lines 6:10). We assume that both $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are non-empty and that $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into bins $\{B_{1},\dots,B_{m}\}$. Each owner then generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{1\leq\ell\leq m}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ represents an independently generated MF statistic for the join key attribute $A_{\mathsf{j}}$, calculated over a specific subset of data filtered by the attribute $A_{\mathsf{ft}}$, such that
@@
here $\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}$ is a group-by-count operation over $A_{\mathsf{j}}$, and $\widehat{\max}_{\epsilon}$ is a report noisy max mechanism (Dwork et al., 2014). It first adds i.i.d. noise from the exponential distribution $\mathsf{Exp}(\frac{2}{\epsilon})$ to each grouped count, then outputs the largest noisy count. We stress that $\mathcal{M}_{\mathtt{synop}}$ will not generate noisy MFs for non-join key attributes, and when $A_{\mathsf{ft}}$ is empty, a global MF will be generated instead of MF tables (Alg 1:10). Moreover, since SPECIAL enables owners to label attributes as unique-valued, if $A_{\mathsf{j}}$ is known to be unique-valued, then $\smash{\widehat{\mathtt{mf}}(A_{\mathsf{j}},\cdot)}$ is always 1. Nevertheless, as exponential noises are non-negative, thus $\widehat{\mathtt{mf}}\geq\mathtt{mf}$ holds for all cases.
@@
[…]
For space concern, we move complete proofs to the full version (Chenghong Wang, 2024). In a sketch, adding $\mathsf{Lap}^{+}$ (or $\mathsf{Lap}^{-}$) to a single bin is $(\epsilon,\delta)$-DP. By parallel and sequential composition, generating $\mathbf{{H}}$ is $(2\epsilon,2\delta)$-DP. Moreover, each noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP. In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,2\delta)$-DP. Given there are in total $c$ such pairs, and thus the total privacy loss is subject to $c$-fold advanced composition (Dwork et al., 2014).
Synopsis transformations. We say that one can perform transformations on released synopses without incurring extra privacy loss, per the post-processing theorem of DP (Dwork et al., 2014). Now, we outline the key synopsis transformation relevant to SPECIAL’s design. First, given any (2d) bounding histogram $\mathbf{H}(\mathsf{pair},D)$ with both $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}$ are non-empty, one can derive the (1d) bounding histograms, i.e. $\mathbf{H}(A_{\mathsf{j}},D)$ and $\mathbf{H}(A_{\mathsf{ft}},D)$, for any single attribute $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ by marginal sums $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over $A_{\mathsf{j}}$ or $A_{\mathsf{ft}}$, respectively. This enables the creation of statistics on individual attributes, even when $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are not included as a standalone synopsis attribute. Moreover, it’s possible to derive relevant join key statistics following a selection on the base relation. For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with valsvals. Beyond bounding histograms, join key MFs over pre-filtered data can also be computed by